OpenStudy (anonymous):

15-8(x-6)=-2.25 what does x equal

OpenStudy (hba):

a(b-c)=ab-ac

OpenStudy (anonymous):

use distributive property... 5-8x-48=-2.25 can you do it from ther?

OpenStudy (anonymous):

Yes I Can